OpenStudy (loser66):

Let {X_n} be a sequence of real numbers such that \(lim_{n\rightarrow \infty} X_n =0\). Show that for any bounded sequence{Y_n}, \(lim_{n\rightarrow\infty}X_nY_n =0\) Please, help.

OpenStudy (loser66):

I have to use Squeeze Theorem to show while I can prove it simpler. By Limit operation theorem, \(lim_{n--> \infty}X_nY_n = lim X_n*lim Y_n= 0* y=0\) But I don't know how to apply Squeeze

OpenStudy (anonymous):

You can't use that limit operation though. You dont know if the \(\lim_{n\rightarrow \infty}y_n\) exists or not.

OpenStudy (anonymous):

Since the sequence \(y_n\) is bounded, we know there exists an \(M>0\) such that : $$0\le\left|y_n\right|\le M$$for all \(n\in \mathbb{N}\). Therefore: $$0\le \left|x_ny_n\right|=\left|x_n\right|\left|y_n\right|\le \left|x_n\right|\cdot M.$$Try using the squeeze theorem here.

OpenStudy (mathmath333):

as it is a bounded sequence u can use ,\(\large\tt \color{black}{Y_n=m_n\times X_n}\) where m is some constant

OpenStudy (loser66):

I am sorry for being late; my computer is stupid

OpenStudy (anonymous):

No worries, I'm running around cleaning. Company coming over later.

OpenStudy (loser66):

Thank you so much.